# Assessing Small Language Models for Code Generation: An Empirical Study with Benchmarks

### v3

@@ 2.6 Data Analysis
The analysis of results in this study involved a combination of descriptive statistical techniques (e.g., mean, median, and standard deviation), hypothesis testing, and comparative ranking [15]. To effectively interpret the experimental outcomes, we utilised a range of visualisations, including bar charts, line graphs, box plots, and bubble plots, to illustrate the various results obtained.
@@
This score integrates both variability measures into a unified indicator of performance stability. To ensure comparability across models, the resulting scores were normalised between 0 and 1 using min-max normalisation, where a higher score indicates greater stability and consistency in performance. For visualisation, we used grouped bar plots to compare model rankings based on mean pass@k and Stability Score across different k-values, and line plots to illustrate trends in SD and CV for each model. Additionally, we statistically assessed whether model performance varied significantly across different groups and benchmarks by conducting a two-way ANOVA test [16]. This two-way ANOVA test evaluated the influence of groups and benchmarks on the observed pass@k scores, examining both main effects and their interaction. Following the ANOVA test, Tukey’s post-hoc test [17] was applied to identify statistically significant differences among the tested conditions. This post hoc analysis aligns best with our experiment, as it supports systematic pairwise comparisons in settings with relatively balanced groups, such as our evaluation setup. Together, the evaluation metrics and statistical tests provided a thorough approach to analysing model performance, enabling us to assess both accuracy and consistency.
@@
To answer RQ2, we analysed the computational efficiency of SLMs in terms of their resource usage and inference speed. We used the same three benchmarks as in RQ1. For each model, we computed the VRAM usage and inference time across the benchmarks. VRAM usage and inference time are widely recognised as key indicators of computational efficiency, as they directly reflect the hardware resource demands and latency of model response [18]. Then we aggregate these scores by calculating their mean to understand the models overall efficiency. Besides, we statistically evaluate the models computational efficiency across different groups by conducting a one-way ANOVA test, because our analysis involved a single independent variable and its effect on computational efficiency metrics. This test evaluates the effect of groups on the resource consumption (VRAM usage) and inference speed. Following this, we applied Tukey’s post-hoc test to compare group-wise differences in resource usage. Additionally, to examine the trade-offs between model performance and efficiency, we visualised the aggregated results using a bubble chart.
@@
To address RQ3, we evaluated SLMs multilingual capabilities using two benchmarks. These benchmarks are explicitly designed to assess code generation capabilities across multiple programming languages. In contrast, the benchmarks we used in RQ1 and RQ2 (HumanEval, MBPP, and Mercury) primarily target only on Python and do not support multilingual evaluation. HumanEvalPack includes three different types of tasks to assess code generation capabilities in four programming languages using the pass@k metric. To obtain a single performance measure per language, we aggregated the scores by computing the mean pass@k across tasks. We then conducted a one-way ANOVA test, as our analysis involved a single independent variable and its effect on performance. The test results evaluate whether the observed differences in performance across languages were statistically significant or not. To complement this analysis, we used box plots to visualise the distribution of model performance by language in different mean pass@k levels, capturing variation and outliers that indicate consistency or disparity in language capabilities.
@@ 3.1 SLMs Performance (RQ1)
| Model | Parameters | HumanEval | MBPP | Mercury | MEAN | ||||||||||||
| pass@1 | pass@2 | pass@5 | pass@10 | pass@1 | pass@2 | pass@5 | pass@10 | pass@1 | pass@2 | pass@5 | pass@10 | pass@1 | pass@2 | pass@5 | pass@10 | ||
| Group 1: SLMs parameter size 1.5B | |||||||||||||||||
| PolyCoder_0.4B | 0.4B | 0.03 | 0.04 | 0.04 | 0.04 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 0.01 | 0.01 | 0.01 | 0.01 | 0.02 | 0.02 | 0.02 |
| Incoder_1.3B | 1.3B | 0.09 | 0.11 | 0.13 | 0.15 | 0.09 | 0.13 | 0.18 | 0.22 | 0.01 | 0.01 | 0.02 | 0.02 | 0.06 | 0.08 | 0.11 | 0.13 |
| Phi1_1.3B | 1.3B | 0.52 | 0.56 | 0.60 | 0.61 | 0.22 | 0.25 | 0.27 | 0.29 | 0.49 | 0.54 | 0.59 | 0.61 | 0.41 | 0.45 | 0.49 | 0.50 |
| DeepSeek-Coder_1.3B | 1.3B | 0.32 | 0.36 | 0.42 | 0.43 | 0.44 | 0.49 | 0.54 | 0.58 | 0.54 | 0.60 | 0.66 | 0.70 | 0.43 | 0.48 | 0.54 | 0.57 |
| OpenCodeInterpreter_1.3B | 1.3B | 0.59 | 0.65 | 0.70 | 0.72 | 0.44 | 0.47 | 0.50 | 0.52 | 0.56 | 0.61 | 0.66 | 0.69 | 0.53 | 0.58 | 0.62 | 0.64 |
| Yi-Coder_1.5B | 1.5B | 0.38 | 0.42 | 0.47 | 0.50 | 0.46 | 0.51 | 0.56 | 0.59 | 0.63 | 0.68 | 0.72 | 0.75 | 0.49 | 0.54 | 0.58 | 0.61 |
| Qwen2.5-Coder_1.5B | 1.5B | 0.44 | 0.51 | 0.58 | 0.61 | 0.51 | 0.58 | 0.64 | 0.68 | 0.66 | 0.71 | 0.74 | 0.76 | 0.54 | 0.60 | 0.65 | 0.68 |
| Group 2: SLMs parameter size 1.5B to 3B | |||||||||||||||||
| CodeGemma_2.0B | 2.0B | 0.07 | 0.09 | 0.12 | 0.15 | 0.31 | 0.40 | 0.48 | 0.52 | 0.47 | 0.53 | 0.58 | 0.60 | 0.28 | 0.34 | 0.39 | 0.42 |
| PolyCoder_2.7B | 2.7B | 0.06 | 0.07 | 0.08 | 0.09 | 0.02 | 0.03 | 0.06 | 0.09 | 0.00 | 0.01 | 0.02 | 0.02 | 0.03 | 0.04 | 0.05 | 0.07 |
| Stable Code_3.0B | 3.0B | 0.21 | 0.25 | 0.29 | 0.31 | 0.25 | 0.31 | 0.38 | 0.43 | 0.36 | 0.47 | 0.60 | 0.65 | 0.27 | 0.34 | 0.42 | 0.46 |
| StarCoder2_3.0B | 3.0B | 0.32 | 0.36 | 0.42 | 0.47 | 0.41 | 0.47 | 0.53 | 0.56 | 0.54 | 0.60 | 0.64 | 0.67 | 0.42 | 0.48 | 0.53 | 0.57 |
| Qwen2.5-Coder_3.0B | 3.0B | 0.51 | 0.57 | 0.63 | 0.66 | 0.57 | 0.63 | 0.68 | 0.71 | 0.68 | 0.72 | 0.76 | 0.78 | 0.59 | 0.64 | 0.69 | 0.72 |
| Group 3: SLMs parameter size 3B to 10B | |||||||||||||||||
| Incoder_6.7B | 6.7B | 0.15 | 0.18 | 0.22 | 0.24 | 0.15 | 0.21 | 0.27 | 0.31 | 0.01 | 0.01 | 0.01 | 0.02 | 0.10 | 0.13 | 0.17 | 0.19 |
| DeepSeek-Coder_6.7B | 6.7B | 0.46 | 0.52 | 0.60 | 0.66 | 0.56 | 0.62 | 0.67 | 0.69 | 0.67 | 0.71 | 0.74 | 0.76 | 0.56 | 0.62 | 0.67 | 0.70 |
| OpenCodeInterpreter_6.7B | 6.7B | 0.73 | 0.76 | 0.79 | 0.80 | 0.59 | 0.63 | 0.66 | 0.67 | 0.70 | 0.72 | 0.74 | 0.75 | 0.67 | 0.70 | 0.73 | 0.74 |
| CodeLlama_7.0B | 7.0B | 0.30 | 0.34 | 0.39 | 0.42 | 0.37 | 0.44 | 0.49 | 0.53 | 0.45 | 0.54 | 0.64 | 0.71 | 0.37 | 0.44 | 0.51 | 0.55 |
| StarCoder2_7.0B | 7.0B | 0.35 | 0.40 | 0.46 | 0.50 | 0.45 | 0.51 | 0.56 | 0.59 | 0.60 | 0.65 | 0.69 | 0.71 | 0.47 | 0.52 | 0.57 | 0.60 |
| CodeGemma_7.0B | 7.0B | 0.42 | 0.52 | 0.65 | 0.74 | 0.46 | 0.57 | 0.67 | 0.72 | 0.60 | 0.68 | 0.73 | 0.75 | 0.49 | 0.59 | 0.68 | 0.74 |
| Qwen2.5-Coder_7.0B | 7.0B | 0.62 | 0.69 | 0.74 | 0.77 | 0.63 | 0.68 | 0.73 | 0.76 | 0.70 | 0.73 | 0.75 | 0.77 | 0.65 | 0.70 | 0.74 | 0.77 |
| Yi-Coder_9.0B | 9.0B | 0.51 | 0.58 | 0.66 | 0.70 | 0.57 | 0.64 | 0.70 | 0.73 | 0.69 | 0.73 | 0.77 | 0.77 | 0.59 | 0.65 | 0.71 | 0.73 |
| Model name | Parameter | pass@1 | pass@2 | pass@5 | pass@10 | ||||||||
| Standard Deviation | Coefficient of Variation | Stability Score | Standard Deviation | Coefficient of Variation | Stability Score | Standard Deviation | Coefficient of Variation | Stability Score | Standard Deviation | Coefficient of Variation | Stability Score | ||
| Group 1: SLMs parameter size 1.5B | |||||||||||||
| PolyCoder_0.4B | 0.4B | 0.02 | 1.73 | 0.00 | 0.02 | 1.25 | 0.02 | 0.02 | 0.87 | 0.03 | 0.02 | 0.65 | 0.02 |
| Incoder_1.3B | 1.3B | 0.05 | 0.73 | 0.17 | 0.06 | 0.77 | 0.06 | 0.08 | 0.74 | 0.02 | 0.10 | 0.78 | 0.01 |
| Phi1_1.3B | 1.3B | 0.17 | 0.40 | 0.09 | 0.17 | 0.39 | 0.05 | 0.19 | 0.39 | 0.01 | 0.18 | 0.37 | 0.01 |
| DeepSeek-Coder_1.3B | 1.3B | 0.11 | 0.25 | 0.23 | 0.12 | 0.25 | 0.12 | 0.12 | 0.22 | 0.04 | 0.14 | 0.24 | 0.02 |
| OpenCodeInterpreter_1.3B | 1.3B | 0.08 | 0.15 | 0.43 | 0.09 | 0.16 | 0.19 | 0.11 | 0.17 | 0.06 | 0.11 | 0.17 | 0.03 |
| Yi-Coder_1.5B | 1.5B | 0.13 | 0.26 | 0.19 | 0.13 | 0.25 | 0.10 | 0.13 | 0.22 | 0.04 | 0.13 | 0.21 | 0.02 |
| Qwen2.5-Coder_1.5B | 1.5B | 0.11 | 0.21 | 0.25 | 0.10 | 0.17 | 0.17 | 0.08 | 0.12 | 0.09 | 0.08 | 0.11 | 0.06 |
| Group 2: SLMs parameter size 1.5B to 3B | |||||||||||||
| CodeGemma_2.0B | 2.0B | 0.20 | 0.71 | 0.01 | 0.23 | 0.66 | 0.00 | 0.24 | 0.62 | 0.00 | 0.24 | 0.57 | 0.00 |
| PolyCoder_2.7B | 2.7B | 0.03 | 1.15 | 0.07 | 0.03 | 0.83 | 0.07 | 0.03 | 0.57 | 0.05 | 0.04 | 0.61 | 0.02 |
| Stable Code_3.0B | 3.0B | 0.08 | 0.28 | 0.32 | 0.11 | 0.33 | 0.10 | 0.16 | 0.38 | 0.02 | 0.17 | 0.37 | 0.01 |
| StarCoder2_3.0B | 3.0B | 0.11 | 0.26 | 0.23 | 0.12 | 0.25 | 0.12 | 0.11 | 0.21 | 0.05 | 0.10 | 0.18 | 0.03 |
| Qwen2.5-Coder_3.0B | 3.0B | 0.09 | 0.15 | 0.40 | 0.08 | 0.12 | 0.28 | 0.07 | 0.10 | 0.12 | 0.06 | 0.08 | 0.08 |
| Group 3: SLMs parameter size 3B to 10B | |||||||||||||
| Incoder_6.7B | 6.7B | 0.08 | 0.78 | 0.10 | 0.11 | 0.81 | 0.03 | 0.14 | 0.83 | 0.01 | 0.15 | 0.80 | 0.00 |
| DeepSeek-Coder_6.7B | 6.7B | 0.11 | 0.19 | 0.29 | 0.10 | 0.15 | 0.19 | 0.07 | 0.10 | 0.11 | 0.05 | 0.07 | 0.09 |
| OpenCodeInterpreter_6.7B | 6.7B | 0.07 | 0.11 | 0.52 | 0.07 | 0.09 | 0.34 | 0.07 | 0.09 | 0.12 | 0.07 | 0.09 | 0.07 |
| CodeLlama_7.0B | 7.0B | 0.08 | 0.20 | 0.40 | 0.10 | 0.23 | 0.15 | 0.13 | 0.25 | 0.04 | 0.15 | 0.26 | 0.02 |
| StarCoder2_7.0B | 7.0B | 0.13 | 0.27 | 0.19 | 0.13 | 0.24 | 0.11 | 0.12 | 0.20 | 0.05 | 0.11 | 0.18 | 0.03 |
| CodeGemma_7.0B | 7.0B | 0.09 | 0.19 | 0.32 | 0.08 | 0.14 | 0.24 | 0.04 | 0.06 | 0.21 | 0.02 | 0.02 | 0.36 |
| Qwen2.5-Coder_7.0B | 7.0B | 0.04 | 0.07 | 1.00 | 0.03 | 0.04 | 1.00 | 0.01 | 0.01 | 1.00 | 0.01 | 0.01 | 1.00 |
| Yi-Coder_9.0B | 9.0B | 0.09 | 0.16 | 0.37 | 0.08 | 0.12 | 0.28 | 0.06 | 0.08 | 0.15 | 0.04 | 0.05 | 0.15 |
@@
We evaluate the SLMs performance on two main aspects: (i) functional correctness of their generated outputs and (ii) stability of generating functionally correct solutions across three benchmarks, which are HumanEval, MBPP, and Mercury. Functional correctness is measured using the pass@k metric. This metric indicates the likelihood that at least one of the top-k generated outputs passes all the test cases. It is well-suited for evaluating code generation models, as it reflects their ability to produce correct outputs in stochastic generation scenarios. Table 3 reports the pass@k scores (k = 1, 2, 5, 10) for each model on these benchmarks, along with their mean scores. These results reflect the functional correctness of the models and provide a basis for comparing their overall performance. Figure 2 illustrates the rank@k positions of the models based on their mean pass@k scores. This ranking reveals several noteworthy pieces of information about model performance across different pass@k levels. Generally, models from Group 3 tend to occupy the top positions in the rankings. For instance, OpenCodeInterpreter 6.7B, Qwen2.5-Coder 7B, and Yi-Coder 9.0B consistently rank among the top three across all pass@k values, with mean pass@1 scores ranging from 0.59 to 0.67. However, some smaller models also achieve high rankings. Qwen2.5-Coder 3.0B from Group 2 consistently secures a position in the top five, achieving a mean pass@1 of 0.59. In addition, some Group 1 models, such as OpenCodeInterpreter 1.3B and Qwen2.5-Coder 1.5B, rank within the top eight, outperforming several Group 3 models, including CodeLlama 7.0B and StarCoder2 7.0B. In contrast, models such as PolyCoder (0.4B and 2.7B) and Incoder (1.3B and 6.7B), from Groups 1, 2, and 3, consistently appear at the lower end of the rankings, with pass@1 scores below 0.10. These trends remain consistent across pass@2, pass@5, and pass@10 levels.

### v4

@@ 2.6 Data Analysis
The analysis of results in this study involved a combination of descriptive statistical techniques (e.g., mean, median, and standard deviation), hypothesis testing, and comparative ranking [inbook]. To effectively interpret the experimental outcomes, we utilised a range of visualisations, including line graphs, box plots, and bubble plots, to illustrate the various results obtained.
@@
This score integrates both variability measures into a unified indicator of performance stability. To ensure comparability across models, the resulting scores were normalised between 0 and 1 using min-max normalisation, where a higher score indicates greater stability and consistency in performance. For visualisation, we used line plots to illustrate trends in SD and CV for each model in different pass@k. Additionally, we statistically assessed whether model performance varied significantly across different groups and benchmarks by conducting a two-way ANOVA test [book]. This two-way ANOVA test evaluated the influence of groups and benchmarks on the observed pass@k scores, examining both main effects and their interaction. Following the ANOVA test, Tukey’s post-hoc test [Tukey] was applied to identify statistically significant differences among the tested conditions. This post hoc analysis aligns best with our experiment, as it supports systematic pairwise comparisons in settings with relatively balanced groups, such as our evaluation setup. Together, the evaluation metrics and statistical tests provided a thorough approach to analysing model performance, enabling us to assess both accuracy and consistency.
@@
To answer RQ2, we analysed the computational efficiency of SLMs in terms of their resource usage and inference speed. We used the same three benchmarks as in RQ1. For each model, we computed the VRAM usage and inference time across the benchmarks. VRAM usage and inference time are widely recognised as key indicators of computational efficiency, as they directly reflect the hardware resource demands and latency of model response [DistilBERT]. Then we aggregate these scores by calculating their mean to understand the models overall efficiency. Besides, we statistically evaluate the models computational efficiency across different groups by conducting a one-way ANOVA test, because our analysis involved a single independent variable and its effect on computational efficiency metrics. This test evaluates the effect of groups on the resource consumption (VRAM usage) and inference speed. Following this, we applied Tukey’s post-hoc test to compare group-wise differences in resource usage. Additionally, to examine the trade-offs between model performance and efficiency, we visualised the aggregated results using a bubble chart.
@@
To address RQ3, we evaluated SLMs multilingual capabilities using two benchmarks explicitly designed to assess code generation capabilities across multiple programming languages. In contrast, the benchmarks we used in RQ1 and RQ2 (HumanEval, MBPP, and Mercury) primarily target only on Python and do not support multilingual evaluation. HumanEvalPack includes three different types of tasks to assess code generation capabilities in four programming languages using the pass@k metric. To obtain a single performance measure per language, we aggregated the scores by computing the mean pass@k across tasks. We then conducted a one-way ANOVA test, as our analysis involved a single independent variable and its effect on performance. The test results evaluate whether the observed differences in performance across languages were statistically significant or not. To complement this analysis, we used box plots to visualise the distribution of model performance by language in different mean pass@k levels, capturing variation and outliers that indicate consistency or disparity in language capabilities.
@@ 3.1 SLMs Performance (RQ1)
Model_Parameter HumanEval MBPP Mercury MEAN pass@1 pass@2 pass@5 pass@10 pass@1 pass@2 pass@5 pass@10 pass@1 pass@2 pass@5 pass@10 pass@1 pass@2 pass@5 pass@10 Group 1: SLMs parameter size 1.5B PolyCoder_0.4B 0.03 0.04 0.04 0.04 0.00 0.00 0.01 0.01 0.00 0.01 0.01 0.01 0.01 0.02 0.02 0.02 Incoder_1.3B 0.09 0.11 0.13 0.15 0.09 0.13 0.18 0.22 0.01 0.01 0.02 0.02 0.06 0.08 0.11 0.13 Phi1_1.3B 0.52 0.56 0.60 0.61 0.22 0.25 0.27 0.29 0.49 0.54 0.59 0.61 0.41 0.45 0.49 0.50 DeepSeek-Coder_1.3B 0.32 0.36 0.42 0.43 0.44 0.49 0.54 0.58 0.54 0.60 0.66 0.70 0.43 0.48 0.54 0.57 OpenCodeInterpreter_1.3B 0.59 0.65 0.70 0.72 0.44 0.47 0.50 0.52 0.56 0.61 0.66 0.69 0.53 0.58 0.62 0.64 Yi-Coder_1.5B 0.38 0.42 0.47 0.50 0.46 0.51 0.56 0.59 0.63 0.68 0.72 0.75 0.49 0.54 0.58 0.61 Qwen2.5-Coder_1.5B 0.44 0.51 0.58 0.61 0.51 0.58 0.64 0.68 0.66 0.71 0.74 0.76 0.54 0.60 0.65 0.68 Group 2: SLMs parameter size 1.5B to 3B CodeGemma_2.0B 0.07 0.09 0.12 0.15 0.31 0.40 0.48 0.52 0.47 0.53 0.58 0.60 0.28 0.34 0.39 0.42 PolyCoder_2.7B 0.06 0.07 0.08 0.09 0.02 0.03 0.06 0.09 0.00 0.01 0.02 0.02 0.03 0.04 0.05 0.07 Stable Code_3.0B 0.21 0.25 0.29 0.31 0.25 0.31 0.38 0.43 0.36 0.47 0.60 0.65 0.27 0.34 0.42 0.46 StarCoder2_3.0B 0.32 0.36 0.42 0.47 0.41 0.47 0.53 0.56 0.54 0.60 0.64 0.67 0.42 0.48 0.53 0.57 Qwen2.5-Coder_3.0B 0.51 0.57 0.63 0.66 0.57 0.63 0.68 0.71 0.68 0.72 0.76 0.78 0.59 0.64 0.69 0.72 Group 3: SLMs parameter size 3B to 10B Incoder_6.7B 0.15 0.18 0.22 0.24 0.15 0.21 0.27 0.31 0.01 0.01 0.01 0.02 0.10 0.13 0.17 0.19 DeepSeek-Coder_6.7B 0.46 0.52 0.60 0.66 0.56 0.62 0.67 0.69 0.67 0.71 0.74 0.76 0.56 0.62 0.67 0.70 OpenCodeInterpreter_6.7B 0.73 0.76 0.79 0.80 0.59 0.63 0.66 0.67 0.70 0.72 0.74 0.75 0.67 0.70 0.73 0.74 CodeLlama_7.0B 0.30 0.34 0.39 0.42 0.37 0.44 0.49 0.53 0.45 0.54 0.64 0.71 0.37 0.44 0.51 0.55 StarCoder2_7.0B 0.35 0.40 0.46 0.50 0.45 0.51 0.56 0.59 0.60 0.65 0.69 0.71 0.47 0.52 0.57 0.60 CodeGemma_7.0B 0.42 0.52 0.65 0.74 0.46 0.57 0.67 0.72 0.60 0.68 0.73 0.75 0.49 0.59 0.68 0.74 Qwen2.5-Coder_7.0B 0.62 0.69 0.74 0.77 0.63 0.68 0.73 0.76 0.70 0.73 0.75 0.77 0.65 0.70 0.74 0.77 Yi-Coder_9.0B 0.51 0.58 0.66 0.70 0.57 0.64 0.70 0.73 0.69 0.73 0.77 0.77 0.59 0.65 0.71 0.73
| Model_Parameter | pass@1 | pass@2 | pass@5 | pass@10 |
| Group 1: SLMs parameter size 1.5B | ||||
| PolyCoder_0.4B | 20 | 20 | 20 | 20 |
| Incoder_1.3B | 18 | 18 | 18 | 18 |
| Phi1_1.3B | 13 | 13 | 14 | 14 |
| DeepSeek-Coder_1.3B | 11 | 11 | 11 | 11 |
| OpenCodeInterpreter_1.3B | 7 | 8 | 8 | 8 |
| Yi-Coder_1.5B | 9 | 9 | 9 | 9 |
| Qwen2.5-Coder_1.5B | 6 | 6 | 7 | 7 |
| Group 2: SLMs parameter size 1.5B to 3B | ||||
| CodeGemma_2.0B | 15 | 16 | 16 | 16 |
| PolyCoder_2.7B | 19 | 19 | 19 | 19 |
| Stable Code_3.0B | 16 | 15 | 15 | 15 |
| StarCoder2_3.0B | 12 | 12 | 12 | 12 |
| Qwen2.5-Coder_3.0B | 4 | 4 | 4 | 5 |
| Group 3: SLMs parameter size 3B to 10B | ||||
| Incoder_6.7B | 17 | 17 | 17 | 17 |
| DeepSeek-Coder_6.7B | 5 | 5 | 6 | 6 |
| OpenCodeInterpreter_6.7B | 1 | 1 | 2 | 2 |
| CodeLlama_7.0B | 14 | 14 | 13 | 13 |
| StarCoder2_7.0B | 10 | 10 | 10 | 10 |
| CodeGemma_7.0B | 8 | 7 | 5 | 3 |
| Qwen2.5-Coder_7.0B | 2 | 2 | 1 | 1 |
| Yi-Coder_9.0B | 3 | 3 | 3 | 4 |
@@
We evaluate the SLMs performance on two main aspects: (i) functional correctness of their generated outputs and (ii) stability of generating functionally correct solutions across three benchmarks, which are HumanEval, MBPP, and Mercury. Functional correctness is measured using the pass@k metric. This metric indicates the likelihood that at least one of the top-k generated outputs passes all the test cases. It is well-suited for evaluating code generation models, as it reflects their ability to produce correct outputs in stochastic generation scenarios. Table 4 reports the pass@k scores (k = 1, 2, 5, 10) for each model on these benchmarks, along with their mean scores. These results reflect the functional correctness of the models and provide a basis for comparing their overall performance. Besides, Table 5 presents the rank@k positions of the models based on their mean pass@k scores. This ranking reveals several noteworthy pieces of information about model performance across different pass@k levels.
Across the benchmarks, models from Group 3 where parameter size 3B and 10B, generally achieve the highest pass@k scores and occupy the top positions in the rankings. For instance, OpenCodeInterpreter 6.7B, Qwen2.5-Coder 7B, and Yi-Coder 9.0B consistently rank among the top four across all pass@k values, with mean pass@1 scores ranging from 0.59 to 0.67, outperforming the smaller SLMs from Groups 1 and 2. This trend persists across pass@2, pass@5, and pass@10, reflecting the advantage of larger model capacity in producing correct code samples. However, the results also show that several mid-sized and small models perform competitively, particularly within Group 1 and Group 2. Qwen2.5-Coder 3.0B from Group 2 achieves a mean pass@1 score of 0.59, placing it among the top five models and performing comparably to larger Group 3 models. Similarly, Group 1 models such as OpenCodeInterpreter 1.3B and Qwen2.5-Coder 1.5B rank within the top eight with a mean pass@1 score of 0.53 and 0.54, outperforming several larger SLMs (e.g., CodeLlama 7B and StarCoder2 7B with a mean pass@1 score of 0.37 and 0.47).
| Model_Parameter | pass@1 | pass@2 | pass@5 | pass@10 | ||||||||
| SD | COV | NSS | SD | COV | NSS | SD | COV | NSS | SD | COV | NSS | |
| Group 1: SLMs parameter size 1.5B | ||||||||||||
| PolyCoder_0.4B | 0.02 | 1.73 | 0.00 | 0.02 | 1.25 | 0.02 | 0.02 | 0.87 | 0.03 | 0.02 | 0.65 | 0.02 |
| Incoder_1.3B | 0.05 | 0.73 | 0.17 | 0.06 | 0.77 | 0.06 | 0.08 | 0.74 | 0.02 | 0.10 | 0.78 | 0.01 |
| Phi1_1.3B | 0.17 | 0.40 | 0.09 | 0.17 | 0.39 | 0.05 | 0.19 | 0.39 | 0.01 | 0.18 | 0.37 | 0.01 |
| DeepSeek-Coder_1.3B | 0.11 | 0.25 | 0.23 | 0.12 | 0.25 | 0.12 | 0.12 | 0.22 | 0.04 | 0.14 | 0.24 | 0.02 |
| OpenCodeInterpreter_1.3B | 0.08 | 0.15 | 0.43 | 0.09 | 0.16 | 0.19 | 0.11 | 0.17 | 0.06 | 0.11 | 0.17 | 0.03 |
| Yi-Coder_1.5B | 0.13 | 0.26 | 0.19 | 0.13 | 0.25 | 0.10 | 0.13 | 0.22 | 0.04 | 0.13 | 0.21 | 0.02 |
| Qwen2.5-Coder_1.5B | 0.11 | 0.21 | 0.25 | 0.10 | 0.17 | 0.17 | 0.08 | 0.12 | 0.09 | 0.08 | 0.11 | 0.06 |
| Group 2: SLMs parameter size 1.5B to 3B | ||||||||||||
| CodeGemma_2.0B | 0.20 | 0.71 | 0.01 | 0.23 | 0.66 | 0.00 | 0.24 | 0.62 | 0.00 | 0.24 | 0.57 | 0.00 |
| PolyCoder_2.7B | 0.03 | 1.15 | 0.07 | 0.03 | 0.83 | 0.07 | 0.03 | 0.57 | 0.05 | 0.04 | 0.61 | 0.02 |
| Stable Code_3.0B | 0.08 | 0.28 | 0.32 | 0.11 | 0.33 | 0.10 | 0.16 | 0.38 | 0.02 | 0.17 | 0.37 | 0.01 |
| StarCoder2_3.0B | 0.11 | 0.26 | 0.23 | 0.12 | 0.25 | 0.12 | 0.11 | 0.21 | 0.05 | 0.10 | 0.18 | 0.03 |
| Qwen2.5-Coder_3.0B | 0.09 | 0.15 | 0.40 | 0.08 | 0.12 | 0.28 | 0.07 | 0.10 | 0.12 | 0.06 | 0.08 | 0.08 |
| Group 3: SLMs parameter size 3B to 10B | ||||||||||||
| Incoder_6.7B | 0.08 | 0.78 | 0.10 | 0.11 | 0.81 | 0.03 | 0.14 | 0.83 | 0.01 | 0.15 | 0.80 | 0.00 |
| DeepSeek-Coder_6.7B | 0.11 | 0.19 | 0.29 | 0.10 | 0.15 | 0.19 | 0.07 | 0.10 | 0.11 | 0.05 | 0.07 | 0.09 |
| OpenCodeInterpreter_6.7B | 0.07 | 0.11 | 0.52 | 0.07 | 0.09 | 0.34 | 0.07 | 0.09 | 0.12 | 0.07 | 0.09 | 0.07 |
| CodeLlama_7.0B | 0.08 | 0.20 | 0.40 | 0.10 | 0.23 | 0.15 | 0.13 | 0.25 | 0.04 | 0.15 | 0.26 | 0.02 |
| StarCoder2_7.0B | 0.13 | 0.27 | 0.19 | 0.13 | 0.24 | 0.11 | 0.12 | 0.20 | 0.05 | 0.11 | 0.18 | 0.03 |
| CodeGemma_7.0B | 0.09 | 0.19 | 0.32 | 0.08 | 0.14 | 0.24 | 0.04 | 0.06 | 0.21 | 0.02 | 0.02 | 0.36 |
| Qwen2.5-Coder_7.0B | 0.04 | 0.07 | 1.00 | 0.03 | 0.04 | 1.00 | 0.01 | 0.01 | 1.00 | 0.01 | 0.01 | 1.00 |
| Yi-Coder_9.0B | 0.09 | 0.16 | 0.37 | 0.08 | 0.12 | 0.28 | 0.06 | 0.08 | 0.15 | 0.04 | 0.05 | 0.15 |
Beyond these trends, additional patterns emerge that further highlight the diversity of SLM behaviour. Models such as DeepSeek-Coder 6.7B show consistent performance across all benchmarks, with balanced improvements in pass@1 to pass@10, reflecting robust generalisation capabilities. CodeGemma 7B exhibits substantial gains from pass@1 to pass@10, indicating that it benefits greatly from sampling diversity, as it generates multiple near-correct answer variants. In contrast, models like StableCode 3B and CodeLlama 7B show uneven behaviour across benchmarks. They show relatively weaker performance on HumanEval but noticeably stronger on Mercury, suggesting that some model families may be particularly suited to algorithmic or dataflow-heavy problems rather than natural-language-driven coding tasks. Among smaller models, Phi-1.3B displays strong performance on HumanEval (pass@1 = 0.52) but much weaker results on MBPP (pass@1 = 0.22), likely reflecting differences in training data and domain specialisation. At the lower end, PolyCoder (0.4B and 2.7B) and Incoder (1.3B and 6.7B) show consistently weak performance, with pass@1 values below 0.10 across benchmarks. This pattern remains stable across pass@2, pass@5, and pass@10, indicating that their lower performance is not simply a top-1 effect. These findings indicate that parameter count alone does not fully determine performance; model architecture and training quality also play significant roles.
| Model_Parameter | pass@1 | pass@2 | pass@5 | pass@10 |
| Group 1: SLMs parameter size 1.5B | ||||
| PolyCoder_0.4B | 20 | 19 | 15 | 11 |
| Incoder_1.3B | 15 | 16 | 17 | 18 |
| Phi1_1.3B | 17 | 17 | 18 | 17 |
| DeepSeek-Coder_1.3B | 11 | 10 | 12 | 14 |
| OpenCodeInterpreter_1.3B | 3 | 7 | 8 | 10 |
| Yi-Coder_1.5B | 14 | 13 | 13 | 12 |
| Qwen2.5-Coder_1.5B | 10 | 8 | 7 | 7 |
| Group 2: SLMs parameter size 1.5B to 3B | ||||
| CodeGemma_2.0B | 19 | 20 | 20 | 20 |
| PolyCoder_2.7B | 18 | 15 | 11 | 13 |
| Stable Code_3.0B | 8 | 14 | 16 | 16 |
| StarCoder2_3.0B | 12 | 11 | 9 | 8 |
| Qwen2.5-Coder_3.0B | 5 | 4 | 5 | 5 |
| Group 3: SLMs parameter size 3B to 10B | ||||
| Incoder_6.7B | 16 | 18 | 19 | 19 |
| DeepSeek-Coder_6.7B | 9 | 6 | 6 | 4 |
| OpenCodeInterpreter_6.7B | 2 | 2 | 4 | 6 |
| CodeLlama_7.0B | 4 | 9 | 14 | 15 |
| StarCoder2_7.0B | 13 | 12 | 10 | 9 |
| CodeGemma_7.0B | 7 | 5 | 2 | 2 |
| Qwen2.5-Coder_7.0B | 1 | 1 | 1 | 1 |
| Yi-Coder_9.0B | 6 | 3 | 3 | 3 |